OpenStudy (anonymous):

What is the product written in scientific notation? (2.1 × 10–2)(3.2 × 10–4) A. 6.72 × 108 B. 6.72 × 10–6 C. 5.3 × 102 D. 5.3 × 10–6

OpenStudy (anonymous):

This Is My Semester Test Prt.1

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Multiply the two numbers together. What's \(2.1 \times 3.2\)?

OpenStudy (anonymous):

it is 6.72

OpenStudy (anonymous):

sorry i was telling my mom something

OpenStudy (igreen):

Yep, now add the exponents: \(10^{-2} \times 10^{-4} = 10^{-2+-4}\)

OpenStudy (anonymous):

ok hold on and i will

OpenStudy (anonymous):

do i add 0.000001+-399/100?

OpenStudy (anonymous):

hello?

OpenStudy (igreen):

Huh? You just add -2 + -4

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

ok.. soo what next?

OpenStudy (igreen):

Yep, so that gives us \(10^{-6}\).

OpenStudy (anonymous):

ok

OpenStudy (igreen):

So we have: \(6.72 \times 10^{-6}\)

OpenStudy (anonymous):

can you help with 13 more?

OpenStudy (igreen):

Yikes, that's a lot, I think I should be able to help with some, but not all of them..

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

it wont let me post it

OpenStudy (anonymous):

solve. 4/7 t =4 t=

OpenStudy (anonymous):

is it 7?

OpenStudy (igreen):

Huh?

OpenStudy (anonymous):

4/7 * 7 is 4 right